OpenStudy (nick1232):

3.4+0.7(5-7.9v)=43.3427

OpenStudy (anonymous):

Distribute and coming like terms on the LHS. The rest is just basic arithmetic.

OpenStudy (nick1232):

can u do it ?

OpenStudy (anonymous):

yeah hold on

OpenStudy (anonymous):

3.4+3.5-5.53v=43.3427 6.9-5.53v=43.3427 -5.53v=-6.9+43.3427 -5.53v=36.4427 v=-6.59

OpenStudy (nick1232):

Thanks

OpenStudy (anonymous):

np :)